OpenStudy (hothotleg):

FAN/MEDAL/TESTIMONY The width of a rectangular backyard is 6 yards shorter than the length. The area is 112 square yards. Write an equation that could be used to determine the dimensions.

563blackghost (563blackghost):

It is stated that the width of the rectangular backyard is 6 yards shorter than the length in which we would make an equation out of this....so if we were to subtract the length by 6 then we would get the width so our equation for width is.... \(\Huge{w=l-6}\) Now we need to put this into an equation for area....what is the formula for finding the area of a rectangle?

563blackghost (563blackghost):

@hothotlEG are you there?

OpenStudy (hothotleg):

I just need the answer and explanation, I have to turn this is really quick!

OpenStudy (hothotleg):

@Aveline I don't know if you know about this type of thing?

OpenStudy (mathmale):

Being in a hurry is never very much to your advantage on OpenStudy. Our rules forbid asking for direct answers as well as handing them out. The previous person was right on target asking you for the formula for the area of a rectangle: "we would get the width so our equation for width is.... w=l−6 Now we need to put this into an equation for area....what is the formula for finding the area of a rectangle?"

OpenStudy (hothotleg):

I believe it's A=L*W?

OpenStudy (hothotleg):

@mathmale

OpenStudy (mathmale):

Yes. Thanks for making the effort to find and share this formula. Now, your width is w=l−6 (where l is the length), and obviously your length is just l. To find the area, multiply the length and width together; in other words, mult. (l) and (l-6) together. Equate your algebraic result to the given area, which is 112 sq. yds. Find l. Then, using the formula w=l−6 , find w.

OpenStudy (hothotleg):

I got -5